# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v1

@@ E.2 Evaluation Metrics
E.3 Results
E.3.1 FoR Inherently Bias in LLMs
C-spilt. The zero-shot setting reflects the LLMs’ inherent bias in identifying FoR. Table 10 presents the accuracy for each FoR class in C-split, where sentences explicitly include information about topology and perspectives. We found that some models strongly prefer specific FoR classes. Notably, Gemme2-9B achieves a near-perfect accuracy on external relative FoR but performs poorly on other classes, especially external intrinsic, indicating a notable bias towards external relative. In contrast, GPT4o and Qwen2-72B perform exceptionally in both intrinsic FoR classes. However, they perform poorly in the relative FoRs.
@@ E.3.1 FoR Inherently Bias in LLMs
E.3.2 Behavior with ICL variations
@@
E.3.3 Experiment with different temperatures
@@
Appendix F In-context learning
F.1 FoR Identification
@@
We provide the prompting for each in-context learning. The prompting for zero-shot and few-shot is provided in Listing 1. The instruction answer for these two in-context learning is “Answer only the category without any explanation. The answer should be in the form of {Answer: Category.}"
@@
For the Chain of Thought (CoT), we only modified the instruction answer to “Answer only the category with an explanation. The answer should be in the form of {Explanation: Explanation Answer: Category.}" Similarly to CoT, we only modified the instruction answer to “Answer only the category with an explanation regarding topological, distance, and direction aspects. The answer should be in the form of {Explanation: Explanation Answer: Category.}", respectively. The example responses are provided in Listing 4 for Spatial Guided prompting.
@@
Listing 1: Prompt for finding the frame of reference class of given context.
@@
Listing 2: Spatial expression examples using for few-shots, Chain-of-Thought (CoT), and Spatial Guide (SG) prompting
F.2 Question Answering Example
@@
Listing 3: Spatial expression examples using for few-shots, Chain-of-Thought (CoT), and Spatial Guide (SG) prompting for question-answering.
F.3 Text to Layout
@@
Listing 4: Prompt for generating bounding coordinates to use as the layout for layout-to-image models.

### v2

@@ H.2 Evaluation Metrics
H.3 Results
H.3.1 FoR Inherently Bias in LLMs
C-spilt. The zero-shot setting reflects the LLMs’ inherent bias in identifying FoR. Table 13 presents the accuracy for each FoR class in C-split, where sentences explicitly include information about topology and perspectives. We found that some models strongly prefer specific FoR classes. Notably, Gemme2-9B achieves a near-perfect accuracy on external relative FoR but performs poorly on other classes, especially external intrinsic, indicating a notable bias towards external relative. In contrast, GPT4o and Qwen2-72B perform exceptionally in both intrinsic FoR classes. However, they perform poorly in the relative FoRs. We also
@@ H.3.1 FoR Inherently Bias in LLMs
H.3.2 Behavior with ICL variations
@@
H.3.3 Experiment with different temperatures
@@
Appendix I Study of Human bias in A-spilt
I.1 Experiment Setting
We provide a Google Form with instructions in Listing 1, followed by 150 randomly ordered context-question pairs. These are sampled based on relatum category: 25 questions per case from the A-split, covering four cases (Cow, Car, Box, Pen) for camera-perspective questions, and two cases (Cow, Car) for relatum-perspective questions, as Box and Pen lack intrinsic direction. No time constraints were imposed, and all participants were compensated at the standard research assistant rate. All participants were informed about the study details at the end and asked for consent to use their responses in the analysis. After acquiring the results, we utilize the same metric defined in Section 5.1. Then, we average the accuracy of each participant to derive the human baseline based on this study.
I.2 Results
Individual human results are shown in Table 15. These results indicate that humans rely heavily on their assumptions when interpreting ambiguous frame-of-reference (FoR) contexts. Most participants favored the intrinsic FoR in A-split scenarios, though this preference varied across individuals, suggesting that background or prior visual experience may influence FoR interpretation. This aligns with findings from cognitive studies on how humans describe spatial relations that recently seen examples may influence Coventry et al. (2018). However, our results contrast with other studies suggesting that humans tend to adopt an egocentric (relative) FoR in scenes with fewer landmarks and an allocentric (intrinsic) FoR in those with more landmarks Forster et al. (2023). Since our study does not provide visual input, we cannot infer how participants mentally reconstruct scenes to answer the questions with only two objects in the scene. The next notable point is that while Box and Pen cases appear straightforward for models (Table 1), humans may still struggle with questions in these categories. Humans do not exhibit the same limitations as LLMs in perspective-shifting tasks. When the context and question differ in perspective, humans still achieve over 90% accuracy, whereas LLMs often struggle with this type of reasoning.
Instruction:
You will be provided with a scene description that describes the spatial relationship between two objects.
The scene description may include object orientation, such as A is facing toward the camera, to indicate the direction A is facing.
You will then receive a question asking about the spatial relationship between A and B from either the camera’s perspective or an object’s perspective.
Your task is to answer the question based on your understanding of the given spatial relationship.
Listing 1: Insturction for collecting human results on QA using A-split of FoREST dataset.
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
P #1 $22.7$ $\mathbf{77.3}$ $88.0$ $8.3$ $\mathbf{91.7}$ $96.0$ $92.0$ $96.0$ $93.0$ $20.8$ $\mathbf{79.2}$ $96.0$ $4.2$ $\mathbf{95.8}$ $96.0$ $96.0$
P #2 $8.7$ $\mathbf{91.3}$ $92.0$ $0.0$ $\mathbf{100.0}$ $96.0$ $36.0$ $56.0$ $70.0$ $20.0$ $\mathbf{80.0}$ $100.0$ $16.7$ $\mathbf{83.3}$ $96.0$ $98.0$
P #3 $\mathbf{78.3}$ $21.7$ $92.0$ $\mathbf{75.0}$ $25.0$ $96.0$ $88.0$ $96.0$ $93.0$ $\mathbf{83.3}$ $16.7$ $96.0$ $\mathbf{87.5}$ $12.5$ $96.0$ $96.0$
Human-baseline $36.6$ $\mathbf{63.4}$ $90.7$ $27.8$ $\mathbf{72.2}$ $96.0$ $72.0$ $82.7$ $85.3$ $41.4$ $\mathbf{58.6}$ $97.3$ $36.1$ $\mathbf{63.9}$ $96.0$ $96.7$
Table 15: QA accuracy in the A-Split of human study. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Appendix J Additional Results with GPT-o4-mini-high
We include results from GPT-4o-mini-high, a model configured for enhanced reasoning, for comparison with our previous results in Table 1 and Table 2. In this experiment, the reasoning-effort parameter is set to high. As shown in Table 16, the model exhibits a balanced preference between relative and intrinsic interpretations. Similar to Qwen2VL, it demonstrates strong reasoning ability, achieving high accuracy even when FoR adaptation is required due to its visual reasoning capability. Notably, GPT-o4-mini-high performs well on questions from the relatum perspective, even outperforms the GPT-4o (SG+CoT) in Table 17. However, its performance declines on camera-perspective questions. We believe this is possibly due to training emphasizing alternative perspectives over the camera view, which may reduce their performance from the camera perspective. Therefore, even with a strong reasoning model, the perspective shift remains challenging in the textual domain.
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
o4-mini-high $\mathbf{68.0}$ $31.6$ $92.0$ $\mathbf{69.9}$ $29.7$ $94.7$ $100.0$ $99.4$ $93.4$ $\mathbf{58.5}$ $41.3$ $86.9$ $\mathbf{50.2}$ $49.5$ $93.3$ $87.8$
Human-baseline $36.6$ $\mathbf{63.4}$ $90.7$ $27.8$ $\mathbf{72.2}$ $96.0$ $72.0$ $82.7$ $85.3$ $41.4$ $\mathbf{58.6}$ $97.3$ $36.1$ $\mathbf{63.9}$ $96.0$ $96.7$
Table 16: Additional results of QA accuracy in the A-Split with GPT-o4-mini-high. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Camera perspective Relatum perspective
GPT-4o (0-shot) $79.7$ $45.1$ $39.5$ $90.2$ $64.2$ $46.9$ $88.5$ $98.2$ $34.8$ $67.5$
GPT-4o (CoT) $81.7$ $\mathbf{76.1}$ $\mathbf{82.4}$ $71.5$ $78.8$ $53.0$ $91.1$ $90.6$ $50.8$ $71.9$
GPT-4o (SG+CoT) $\mathbf{97.9}$ $72.2$ $72.7$ $\mathbf{93.4}$ $\mathbf{85.8}$ $48.9$ $96.3$ $95.9$ $36.1$ $71.8$
o4-mini-high $68.7$ $73.1$ $72.1$ $79.0$ $71.4$ $\mathbf{78.9}$ $91.9$ $93.9$ $\mathbf{62.7}$ $\mathbf{84.5}$
Table 17: Additional results of QA accuracy in the C-Split with GPT-o4-mini-high. ER, EI, II, and IR denote external relative, external intrinsic, internal intrinsic, and internal relative FoRs. Avg represents the micro-average accuracy. CP refers to context with camera perspective, while RP denotes context with relatum perspective.
Appendix K LLM Prompt
K.1 FoR Identification
@@ I.2 Results
We provide the prompting for each in-context learning. The prompting for zero-shot and few-shot is provided in Listing 2. The instruction answer for these two in-context learning is “Answer only the category without any explanation. The answer should be in the form of {Answer: Category.}"
@@
For the Chain of Thought (CoT), we only modified the instruction answer to “Answer only the category with an explanation. The answer should be in the form of {Explanation: Explanation Answer: Category.}" Similarly to CoT, we only modified the instruction answer to “Answer only the category with an explanation regarding topological, distance, and direction aspects. The answer should be in the form of {Explanation: Explanation Answer: Category.}", respectively. The example responses are provided in Listing 5 for Spatial Guided prompting.
@@
Listing 2: Prompt for finding the frame of reference class of given context.
@@
Listing 3: Spatial expression examples using for few-shots, Chain-of-Thought (CoT), and Spatial Guide (SG) prompting
K.2 Question Answering Example
@@
Listing 4: Spatial expression examples using for few-shots, Chain-of-Thought (CoT), and Spatial Guide (SG) prompting for question-answering.
K.3 Text to Layout
@@
Listing 5: Prompt for generating bounding coordinates to use as the layout for layout-to-image models.